### v1

# Supervised Fine-Tuning: An Activation Pattern Optimization Process for Attention Heads

Abstract
Though demonstrating promising potential, LLMs’ performance on complex tasks, such as advanced mathematics and complex disease diagnosis is still unsatisfactory. A key issue is the present LLMs learn in a data-driven schema, while the instruction dataset about these complex tasks is both scarce and hard to collect or construct. On the contrary, a prominent phenomenon is that LLMs can learn rather fast on those simpler tasks with adequate prior knowledge captured during pretraining stage. Thus, if the prerequisite and mechanism of such rapid generalization could be elucidated, it could be highly beneficial in enhancing the efficiency and effectiveness of the LLM’s ability to learn complex tasks. Thus, in this paper, we employ a gradient-based method, to dissect the process that the SFT process adapts LLMs to downstream tasks via the perspective of attention patterns. We find that: (1) LLMs selectively activate task-specific attention heads during SFT; (2) activation patterns for complex tasks are combinations of basic task patterns; and (3) changes in a few parameters can significantly impact activation patterns after SFT on a small number of samples. Based on these insights, we conduct experiments to examine whether these conclusions could effectively enhance the efficiency and effectiveness of SFT, particularly in handling complex tasks and when instructional resources are scarce. Our research not only uncovers the underlying reasons behind LLMs’ rapid learning and generalization mechanisms but also provides practical solutions for addressing data challenges in complex and specialized tasks.
Introduction
The Supervised Fine-tuning Process (SFT) is essential for aligning Large Language Models (LLMs) with human objectives to complete downstream tasks (Zhang et al. 2023). Leveraging the extensive prior knowledge acquired during pretraining, fine-tuning LLMs on instruction sets enables them to perform well across various tasks. (Dong et al. 2023; Xia et al. 2024). However, in more complex tasks like mathematical reasoning, LLM performance remains unsatisfactory, indicating there is still considerable room for improvement. (Fourrier et al. 2024; Contributors 2023).
One major reason for such performance limitation is that these complex tasks often demand simultaneous utilization of multiple types of knowledge and skills, and thus, more instructions are required to adapt LLMs to these tasks (Kaplan et al. 2020). However, the more complex the task, the more challenging it is to collect and construct relevant instruction data (Zhang et al. 2023; Minaee et al. 2024). This in turn limits the ability of LLMs to improve performance on these tasks. For instance, data related to complex decision-making processes, such as medical decision-making and social behavior prediction, are often challenging to construct or obtain due to their sensitive nature, proprietary status, or the prohibitively high costs associated with data collection.
Figure 1: Visualization of activation pattern changes in Llama3-8B on the test set before and after SFT with the GSM8K training set.
In contrast, on most basic tasks, LLMs demonstrate prominent rapid generalization (Zhang et al. 2023). By training on just a few thousand instructions, LLMs can learn to complete various tasks(Xia et al. 2024). Therefore, understanding the mechanisms and conditions that enable LLMs’ rapid learning and generalization could strongly guide their adaptation to complex tasks.
To address this issue, we propose to analyze how SFT adapts LLMs to different tasks through the perspective of activation patterns of attention heads using a gradient-based method. Previous studies show that attention heads serve as basic functional units in transformer-based models (Voita et al. 2019; Clark et al. 2019; Hao et al. 2021; Wang et al. 2023). Through pretraining, these attention heads capture different types of information, model various relationships, and demonstrate distinct functions across tasks.
Therefore, analyzing how LLMs utilize different functions to solve tasks during the SFT process could elucidate the mechanisms behind SFT learning and generalization.If an attention head influences outputs in certain tasks, it is heuristically considered “activated”. The composition of activations across different attention heads forms the activation pattern of the LLMs (Figure 1), indicating how the model’s functional units are integrated to solve a specific task.To determine if an attention head influences the output (i.e., is activated), we use a gradient-based method because gradients naturally measure the sensitivity and impact of outputs on input features  (Wang et al. 2024). By calculating the expected value of the gradient on the attention score matrix, we can quantify the impact of each attention head on model performance before and after SFT This approach reveals the changes in LLMs following SFT, and explains how SFT adapts LLMs to downstream tasks.
By analyzing the alternation of attention head activation patterns upon tasks across multiple domains during the SFT process, we find the critical role of the change of activation pattern in fast learning and generalization during the SFT process, as: (1) LLMs learns to complete tasks by selectively activate attention heads during the SFT process. After SFT, the LLMs activates certain tasks-specific attention heads, and the differences in activation patterns between tasks become more pronounced; (2) LLMs complete tasks by learning to integrate basic skills. We observe that the activation patterns on completing complex tasks are a combination of several basic tasks; (3) Changes in a few parameters can significantly impact activation patterns after SFT on a small number of samples. The model’s activation patterns exhibit significant changes compared to the base model in the early stage of the SFT process with rather limited optimization steps, showing that minimal parameter adjustments can significantly alter activation patterns.
Based on these findings, we further validate them by examining whether they can enhance the effectiveness and efficiency of SFT, especially on complex tasks and when instructions are scarce. We conduct two experiments: The first simulates a scenario with limited complex task data. By fine-tuning LLMs using instructions on the basic skills required for these tasks, LLM performance significantly improves. The second experiment addresses scenarios where high-quality domain data is private or unavailable. Based on activation patterns, we select relevant instructions from a large pool to approximate the effects of private data using public instructions, thereby improving LLM performance. These results not only support our findings on rapid learning and generalization mechanisms, but also offer practical solutions for overcoming data challenges in complex and specialized tasks.
Background and Related Work
One factor limiting the effectiveness of LLMs is the opacity of their mechanisms and the lack of understanding of generalization in SFT. Some studies have used gradient-based methods to identify key features and parameters influencing LLM outputs, shedding light on their internal mechanisms (Hao et al. 2021; Wang et al. 2023; Yang, Wang, and Bilgic 2023). However, most research on SFT treats LLMs as black boxes (He et al. 2024; Dong et al. 2023; Lu et al. 2023). Given this black-box approach, there is a gap in understanding the detailed mechanisms behind SFT, which our work aims to address. Our work takes a pioneering approach in this area by exploring these underlying mechanisms.
A key question is which perspective is most effective for this analysis. Previous studies (Yu et al. 2024; Fu et al. 2023) suggest that parameter changes after SFT are limited, and given the vast number of parameters, interpreting their impact is challenging.
In contrast, previous studies have identified that an attention head can act as a basic functional unit in transformer-based models (Voita et al. 2019; Clark et al. 2019; Wang et al. 2023), with different attention heads serving distinct functions. For example, Voita et al. (2019) and Clark et al. (2019) found that attention heads at different layers focus on different parts-of-speech, while Wang et al. (2023) observed that during in-context learning, attention heads at different layers focus on different parts of the in-context learning process. Additionally, prior research (Jin et al. 2024; Lee and Hwang 2024) has suggested that attention heads exhibit task specificity, as pruning certain attention heads does not lead to performance degradation on downstream tasks. Heuristically, since a task can be viewed as a combination of several basic tasks, completing it requires the coordination of multiple corresponding basic functional units, i.e., attention heads. Thus, in this paper, we investigate the fast generalization mechanism of SFT from the perspective of activation patterns in attention heads, using a gradient-based method for our analyses..
For clarity, we define attention heads that impact downstream tasks as activated and those that do not as inactive. The matrix representing the influence of attention heads on downstream tasks is called the activation pattern. By examining how attention head activations change in response to fine-tuning, we aim to uncover the mechanisms enabling rapid learning and generalization in LLMs.
Methodology
During pretraining, LLMs learn to use different attention heads to capture various types of information (Voita et al. 2019; Khaki and Plataniotis 2024). Heuristically, a complex task can often be viewed as a composition of several basic tasks. For instance, solving a mathematical problem using code, as in SGSM (Christ, Kropko, and Hartvigsen 2024), can be broken down into fundamental tasks like code generation and mathematical reasoning. Since each attention head may correspond to a specific function, an LLM can invoke multiple attention heads and integrate their functions to accomplish a complex task. Combining these observations leads to our core assumption: during the SFT process, LLMs quickly adapt to downstream tasks by invoking different attention heads. Thus, in this paper, we investigate the mechanism of fast generalization in the SFT process by analyzing changes in the activation patterns of attention heads.
Gradient-Based Attention Head Activation Patterns Analysis
One critical issue is how to measure if an attention head has been used in a task, in other words, it has been activated. Inspired by (Hao et al. 2021) and (Wang et al. 2023), we measure this using model gradients, as gradients can naturally reflect the subtle effects of input features or parameters on the model’s output (Wang et al. 2024). When the gradient of a particular attention head is large, it indicates that the model’s output is more sensitive to the attention head, suggesting a higher influence on outputs.
Specifically, given an LLMs with $L$ layers and $H$ attention heads per layer, on a dataset $\mathcal{T}$ about one certain task, the activation pattern (AP) could be represented as a $L\times H$ matrix, with each element representing the activation level (AL) reflecting the contribution of one specific attention head to a given task:
$AP^{\mathcal{T}}=\{AL_{l,h}^{\mathcal{T}}\}_{l\in[1,\dots,L],h\in[1,\dots,H]},$ (1)
where each element $AL_{l,h}$ corresponds to the activation level of the $h$-th attention head in the $l$-th layer during the given task, which is measured as:
$AL_{l,h}=\frac{1}{N}\sum_{i}{\Gamma_{l,h}^{T}}\frac{\partial L(x_{i})}{\partial\Gamma_{l,h}},$ (2)
where $N$ represents size $\mathcal{T}$, $x_{i}$ is the $i$th instance in $\mathcal{T}$ and $L(x_{i})$ is the loss value of the model on $x_{i}$. $\Gamma_{l,h}$ is the attention matrix of the $hl$th attention head. $\partial\Gamma_{l,h}$ is the gradient of the $lh$th attention head with respect to $L(x_{i})$. By combining the absolute value of attention scores with the sensitivity of loss on the attention scores, the total influence of a certain attention head on outputs and the task could be quantified. Thus, $AP^{task}$, composed of $AL^{task}$, reflects how each attention head contributes to task completion.
Analytical Framework
To validate our core assumption, we focus on three progressive issues: (1) How do the activation patterns change during the SFT process for each task, and are these changes specific to the task? In other words, can we grasp the characteristics of the task from the perspective of activation patterns? (2) What is the relationship between the activation patterns when learning complex tasks and those when learning basic tasks? This reflects the prerequisites that may be necessary for learning complex tasks; and (3) on basic tasks, how many training samples are required to significantly change the activation pattern? The necessary sample size indicates the model’s ability to achieve rapid generalization.
First, we evaluated changes in activation patterns before and after SFT across tasks. We observed that more attention heads were activated post-SFT, with these activations being task-specific, suggesting that during the SFT Process, Attention Heads are Selectively Activated. Next, we explored the relationship between activation patterns in complex and basic tasks, finding that the patterns of several basic tasks can effectively approximate those of complex tasks. From this, we concluded that activation patterns in complex tasks are combinations of basic task patterns. Finally, by dynamically observing activation pattern changes during SFT, we concluded that even small datasets can significantly alter activation patterns.
To further validate our conclusions, we applied our findings to enhance the effectiveness of SFT on complex tasks. By evaluating the resulting improvements, we further confirmed the validity of our theory.The following sections provide a detailed analysis of the process and the conclusions drawn from the analysis.
SFT Adapts to Downstream Tasks by Modifying Activation Patterns
This section employs gradient-based analysis to examine the impact of SFT on transformer attention head activation patterns. Our analysis produces three key conclusions that address the questions previously posed.
Firstly, after the SFT process, LLMs learn to complete tasks by selectively activating attention heads. Secondly, changes in activation patterns for compound tasks can be interpreted as a combination of changes observed in more basic tasks.Thirdly, during the SFT process, the model’s activation patterns exhibit exponential changes, meaning that even a small amount of data can significantly alter these patterns.
Attention Heads are Selectively Activated
In this section, we first analyze the changes in attention head activation patterns before and after SFT across different tasks. Then, we explore the relationship between these activation pattern changes across tasks.
Analytical Method
To analyze the changes in model activation patterns before and after SFT for each task, we used three metrics: Gini Coefficient, Coefficient of Variation (CV), and Kurtosis. The Gini Coefficient measures distribution inequality, with values above 0.4 indicating significant concentration in a few attention heads. The Coefficient of Variation reflects relative dispersion, with values over 1 indicating substantial variability. Kurtosis describes the distribution’s shape, where values above 3 suggest heavy tails and sharp peaks, indicating the presence of extreme values.
To analyze the relationship between changes in activation patterns across different tasks, we used the Pearson correlation coefficient of activation changes, which reflects the linear consistency of the direction of changes. A high correlation between the activation patterns of two tasks indicates that the attention heads relied upon to complete both tasks are highly similar. Additionally, a high correlation in the rate of change of activation patterns suggests that the degree to which attention heads are engaged during the SFT process is similar for both tasks.
Experiments Setup
We conducted experiments on models with varying performance levels, including Llama3-8B (AI@Meta 2024), Gemma-7B (Team et al. 2024), and OPT-6.7B (Zhang et al. 2022), across a range of tasks: mathematical reasoning (MATH (Hendrycks et al. 2021b), GSM8K (Cobbe et al. 2021)), coding (Code Search Net (Husain et al. 2019), SGSM (Christ, Kropko, and Hartvigsen 2024)), and natural language processing and reasoning (HellaSwag (Zellers et al. 2019), Winogrande (Sakaguchi et al. 2021), ARC (Clark et al. 2018)). Details on the data and training process are provided in Appendix I and II.
Results
Model State Gini CV Kurtosis
Llama3-8B Before SFT 0.50 1.19 95.37
After SFT 0.33 0.71 39.55
Gemma-7B Before SFT 0.48 1.71 240.99
After SFT 0.38 1.23 130.82
OPT-6.7B Before SFT 0.42 1.06 50.67
After SFT 0.38 0.819 24.23
Table 1: Statistics on the distribution of activation patterns for different LLMs. Experiments were conducted on tasks such as Code Search Net, GSM8k, MATH, SGSM, ARC, HellaSwag, and Winogrande.
As shown in Table 1, before SFT, the base models exhibited significant outliers (Kurtosis ¿ 3), uneven distribution (Gini ¿ 0.4), and high variability (CV ¿ 1) in activation patterns across multiple tasks. This suggests that only a few attention heads have high activation levels and are contributing to the task. After the SFT process, the distribution of activation patterns became more uniform, with decreases observed in the Gini, CV, and Kurtosis index. This suggests that SFT adapts LLMs to specific tasks by increasing the activate level of attention heads. However, the distribution of LLMs activation values remains highly skewed, reflecting that only a few heads play a significant role.
Figure 2: The correlation coefficients of the activation pattern change rates for the Llama3-8B, Gemma-7B, and OPT-6.7B models on tasks before SFT, after SFT, and during the SFT process (corresponding to the top, middle, and bottom sections, respectively).
Figure 2 shows the similarity in the activation pattern change rates across different tasks for three models before and after SFT, measured by correlation coefficients. Moreover, as shown in Figure 2, based on the correlation analysis, after SFT, the tasks were divided into two categories: math and code-related tasks, and text reasoning tasks. This largely aligns with human understanding of the relationships between these tasks. Therefore, the attention patterns can, to some extent, reflect the characteristics of the tasks and exhibit task specificity. After SFT, the changes in the model’s activation patterns demonstrate stronger task specificity, as evidenced by the decrease in the correlation coefficients of the attention patterns across different tasks—in other words, the similarity between tasks has decreased.
Considering that the activation levels increase after SFT, we deduce that SFT adapts models to specific tasks by selectively enhancing task-relevant activation patterns.
Activation Patterns on Complex Tasks are Combinations of Basic Tasks
This section focuses on the relationship between activation pattern changes in basic and complex tasks. Understanding this relationship is key to equipping models with the necessary prerequisites for rapid learning of complex tasks.
Analytical Method
Figure 3: The two tasks above are basic tasks that each rely on a single skill, while the one below is a complex task that relies on both coding and mathematical skills.
In this paper, we define complex tasks as those that can be decomposed into more basic skills. For instance, as illustrated in Figure 3, solving the ”chicken and rabbit” problem requires elementary math skills, while constructing a binary tree requires Python coding skills—both considered basic tasks. However, solving the ”chicken and rabbit” problem using Python code is a complex task, as it requires both mathematical reasoning and coding skills.
The actual relationship between activation patterns in basic tasks and complex tasks can be quite intricate. Without generality, in our study, we employ a linear function to model the relationship between the activation patterns of basic and complex tasks. Specifically, we take the changes in activation patterns for complex tasks as the dependent variable, and employ the changes in activation patterns for multiple simple tasks as independent variables, using the least squares method to fit the regression coefficients.
$\Delta AP^{complex}=\sum_{i=1}^{n}\alpha_{i}\Delta AP^{basic_{i}}+\epsilon,$ (3)
where $\alpha_{i}$ is a regression coefficient representing the contribution of each basic task $i$ to the change in activation patterns for the complex task, $n$ is the total number of basic tasks, and $\epsilon$ is the error term. Subsequently, we employ the R-squared ($R^{2}$) value to measure the goodness of fit, which represents the proportion of the variance in the dependent variable that can be explained by the independent variables.
Experiments Setup
In this study, we conducted two groups of experiments: (1) We take MATH, GSM8K, Code Search Net, HellaSwag, Winogrande, and ARC as the basic tasks, and select the SGSM task, which requires coding ability and elementary math skills, as the complex task. (2) We conduct experiments on the Infinity Instruct dataset (of Artificial Intelligence BAAI), which provides each instruction tags describing the skills and knowledge required to complete that instruction, so that we can select both basic and complex task-related instructions. Specifically, we selected instructions requiring only a single skill, including ‘mathematics’, ‘logical reasoning’, ‘programming and software development’, ‘education and consulting’, ‘financial and business knowledge’, and ‘legal knowledge’, to form into the basic task data set. Next, we selected instructions requiring both ‘logical reasoning’ and ‘programming/software development’ skills as the complex tasks for SFT and calculated their activation patterns.
Result
Figure 4: Top: The least squares method was used to fit the activation pattern changes of traditional NLP tasks in SFT to the activation pattern changes in SGSM. A higher $R^{2}$ indicates a better fit, with Code Search Net and GSM8k showing the highest $R^{2}$ values. Bottom: The least squares method was used to fit the activation pattern changes of SFT instruction data to those of tasks requiring both “logical reasoning” skills and “programming and software development” instructions. The combination of “logical reasoning” skills and “programming” instructions achieved the highest $R^{2}$ value.
The upper part of Figure 4 shows the $R^{2}$ values obtained by fitting the activation patterns of various traditional NLP tasks using the least squares method, with the activation pattern changes of the SGSM task as the dependent variable. Among these tasks, Code Search Net and GSM8K yield the highest $R^{2}$ value of 0.97. This indicates that (1) the activation patterns of complex tasks can be well-fitted by the linear combination of activation patterns from simpler tasks; (2) SGSM is a task that involves solving elementary math problems using code, requiring a combination of elementary math and coding skills. In fact, the changes in SGSM’s activation patterns can be well-fitted by the activation pattern changes from the elementary math task GSM8K and the coding task dataset Code Search Net.
A similar phenomenon can be observed in the results based on Infinity Instruct. The lower part of Figure 4 shows the $R^{2}$ values obtained by fitting the activation patterns of various instruction-based NLP tasks using the least squares method, with the activation pattern changes of the instruction tasks that require both logical reasoning and programming/software development skills as the dependent variable. Among these tasks, the instructions that rely solely on logical reasoning and those that rely solely on programming/software development yield the highest $R^{2}$ value of 0.95.
This shows that the activation patterns of complex tasks can be formed by combining the activation patterns of several simpler tasks, and this compositional relationship reflects the real-world connections between these tasks. Therefore, SFT adapts LLMs to complex tasks by integrating the attention heads corresponding to the simpler tasks.
Furthermore, the weigthed sum of activation patterns from models fine-tunned on simple tasks resembles that on complex tasks. It suggest that learned enough from basic tasks, its state will approximate that of a model directly fine-tuned with instructions related to complex tasks. This provides guidance on how to establish the necessary prerequisites for complex task performance.
Small Data Can Change Activation Patterns
In this section, we find that SFT can quickly adapt LLMs to certain basic tasks by adjusting the activation level of attention heads with rather limited instances. This suggests that with sufficient prior knowledge, the model can quickly adapt to a task even with a small amount of data.
Analytical Method
To analyze changes in instruction activation patterns during SFT, we saved checkpoints throughout the process for different tasks. We measured the similarity and distance between the activation patterns at different checkpoints using two metrics: Correlation Coefficient and Mean Squared Error (MSE). Specifically, we calculated the Correlation and MSE between the activation patterns of the model at the $i$-th step of SFT on a given task and the activation patterns of the base model on the same task. The Correlation reflects the consistency in activation pattern changes, while MSE captures the magnitude of these changes.
Experiments and Results
Figure 5: Tracking changes in correlation coefficient and MSE activation patterns of the Llama3-8B, Gemma-7B, and OPT-6.7B models during fine-tuning on datasets including Code Search Net, GSM8k, MATH, SGSM, ARC, HellaSwag, and Winogrande.
As shown in Figure 5, activation patterns change rapidly in the initial steps of SFT. This suggests that fine-tuning with small datasets can significantly reshape a model’s attention patterns and alter its performance. Previous research (Xia et al. 2024) supports this, indicating that rapid learning during SFT is driven by swift changes in attention patterns.
Moreover, the extent of changes in activation patterns correlates with improvements in model performance. As shown in Figure 5, the OPT-6.7B model shows significant changes in activation patterns across most tasks, particularly for complex tasks like MATH. The shifts in activation patterns post-fine-tuning are more pronounced compared to those in the Llama3-8B and Gemma-7B models. This suggests that the initial activation patterns of OPT-6.7B may not be well-suited for solving complex problems like MATH, and thus need more instances to approach the convergence.
By analyzing models with varying capabilities, we find that for simpler tasks, fewer steps, samples, and less SFT data are needed for stronger LLMs to approach convergence and adapt to the task. In contrast, weaker models require more data due to having less prior knowledge from pre-training. This suggests that the foundation of rapid generalization is sufficient prior knowledge. Consequently, with enough prior knowledge, a small number of samples can enable rapid learning of complex tasks. Given that complex tasks can be decomposed into simpler ones, this raises the question: Once a model has sufficient prior knowledge, can it be fine-tuned with just a small number of instructions to rapidly learn complex tasks?
Deduction and Applications
We further validate our conclusions by testing their applicability in enhancing the effectiveness of SFT. Based on our observations of activation patterns, we focus on two scenarios: (1) Can we improve SFT on complex tasks by providing LLMs with prior knowledge of the basic tasks that make up the complex task? (2) Can we identify relevant data from a large candidate pool to approximate the effect of a target dataset, particularly when the target dataset is unavailable?
Scenario 1: Combining Data of Basic Tasks to Promote the Learning of Complex Task
Our analysis indicates that LLMs learn to complete complex tasks by leveraging a combination of their fundamental capabilities. A natural question arises: for a complex task, can we enhance the efficiency and effectiveness of the learning process by equipping LLMs with relevant basic capabilities? We conduct experiments on these settings: (1) Aiming at adapting LLMs to a complex task with a limited number of instructions given; (2) A large number of instructions about various basic tasks are available; (3) The relationship between the complex task and basic tasks are unknown.
Given that the relationships between the complex task and basic tasks are unknown, we first fit these relationships using the regression function described in Equation 3. Intuitively, the regression coefficients $\alpha_{i}$ describe the proportion that the complex task can be “composed ” by a basic task $i$. Thus, we construct a preliminary instruction set by mixing instructions about basic tasks according to the magnitude of the regression coefficients:
$\text{Dataset}^{preliminary}=\{N\times\alpha_{i}/\sum_{i}\alpha_{i}\}_{i=0}^{|B|}$ (4)
where $|B|$ is the number of candidate basic tasks, $N$ is the predefined size of $\text{Dataset}^{preliminary}$. During experiment, LLM is first trained on $\text{Dataset}^{preliminary}$, then finetuned on the target task dataset $\text{Dataset}^{\text{complex task}}$.
Experiment Settings
To simulate scenarios where LLMs lack sufficient preliminary knowledge, we conducted experiments on the Llama series models (Touvron et al. 2023a, b; AI@Meta 2024), which have not been specifically optimized for Chinese corpora and exhibit relatively limited performance on Chinese-related tasks. In this section, we set the target complex task as MathBench (Liu et al. 2024), which requires the model to solve school mathematics questions using Chinese. The MathBench dataset 358 instances, and we use only 100 instances as $\text{Dataset}^{\text{complex task}}$, the left part for testing. In addition to the basic tasks described in Section 4.2, we include RefGPT (Yang et al. 2023) (a dataset of Chinese factual knowledge) as a basic task dataset. After the regression process, we keep only basic datasets with the Top 2 largest regression coefficients, so as to filter out the noises. The remaining basic tasks are GSM8K (English elementary math) and RefGPT.Experiments are conducted on models of similar size but varying capabilities: Llama-7B (Touvron et al. 2023a), Llama2-7B (Touvron et al. 2023b), and Llama3-8B (AI@Meta 2024). Detailed experimental procedures are in Appendix I and II.
Model Base SFT Random Ours
Llama-7B 28.68 31.78 33.82 36.82
Llama2-7B 29.07 36.43 34.51 40.70
Llama3-8B 50.39 52.33 51.16 55.03
Table 2: Performance on the Mathbench test set: ”Base” refers to the base model results, “SFT” refers to results after fine-tuning with 100 Mathbench samples, and “Random” represents the average of five runs trained with a randomly mixed dataset of GSM8K and RefGPT data, followed by fine-tuning with 100 Mathbench samples Best results are bolded, and second-best are underlined.
Analysis
As shown in Table 2, our approach consistently outperforms baseline models across all evaluations, including models fine-tuned solely on the Mathbench training set and those trained with a random mix of elementary mathematics and Chinese instruction data. Compared to models fine-tuned only on Mathbench, our method shows significant improvements in handling complex tasks through subtask training. Additionally, using activation pattern proportions for dataset construction enhances the model’s ability to utilize simple task data for complex tasks, outperforming models trained on randomly combined subtasks. For Llama-7B, learning from randomly proportioned basic task data can still improve complex task performance, but for more powerful models like Llama2-7B and Llama3-8B, this approach may hinder performance. These findings suggest that guiding dataset proportioning with activation patterns from basic tasks significantly enhances the model’s ability to learn and generalize to complex tasks, reinforcing our earlier analysis of the linear relationship between basic and complex tasks.
Scenario 2: Selecting Relevant Data from Candidates
While various open-source Domain LLMs demonstrate strong capabilities, their instruction data is often unavailable. Hence, considering the task-specificity of activation patterns, could we select relevant datasets from open instruction sets and use them to approximate the performance trained upon private domain datasets? Thus, we conduct experiments under the following conditions: (1) only a small amount of developer-constructed pseudo-private data is available;(2) the complete dataset used to train the domain model is unknown;(3) The public data is sufficiently large, and there exists data that can improve the model’s performance on the target task.
Experiment Settings
Model Base T100 R100 T300 R300 T500 R500
Llama-7B 30.00 31.66 30.33 36.15 32.81 39.09 34.68
Llama2-7B 35.08 39.05 37.06 40.51 37.62 40.70 37.95
Llama3-8B 56.75 57.39 56.19 57.95 56.37 57.51 56.19
Table 3: Average accuracy of the model on four tasks: mathematics, physics, chemistry, and biology. ”Base” represents the baseline model’s average accuracy on these tasks. “T100” indicates the accuracy after fine-tuning with 100 most similar data points based on activation patterns. ”R100” represents the accuracy after fine-tuning with 100 randomly selected unrelated MMLU data points, with similar meanings for “T300”, “R300”, “T500”,and “R500”.
To simulate the performance of LLMs when target data is unavailable, we conducted experiments on the MMLU dataset (Hendrycks et al. 2021a), which covers multiple domains. We selected test set data from certain domains as private (target data) and treated the remaining MMLU test set data as non-target data. Additionally, we used validation and dev set data from the same domains as pseudo-private data, simulating the small, manually constructed datasets that developers might create when the original data is unavailable. Fine-tuning on this pseudo-private data enhanced the model’s task specificity for the target domain. We then used Equation 5 to identify the top $m$ non-target data points from MMLU that are most similar to the pseudo-private data, and used these for further training. We compared the performance of this approach with the base model and models trained on an equal amount of randomly selected non-target data. To ensure robustness, we averaged the results across multiple domains, including mathematics, physics, chemistry, and biology, treating them as target data.
$\text{Dataset}^{approx}=\operatorname{Top}_{m}Corr(AP^{target},AP^{D_{i}}),$ (5)
where $\text{Dataset}^{approx}$ represents the entire dataset, $D_{i}$ is the $i$-th sample from $D$, $AP^{D_{i}}$ is the activation pattern for the $i$-th sample after SFT, and $Corr(*,*)$ is the correlation coefficient.
Analysis
As shown in Table 3, models trained with data selected by activation pattern similarity consistently outperform those trained with randomly selected data. This suggests that using the similarity of activation patterns, we can effectively identify data that share similarities with the private data.These experimental results not only support our theoretical hypothesis regarding activation patterns but also demonstrate the broad potential of this strategy in practical applications. By judiciously selecting training data based on activation patterns, we can achieve significant performance improvements in scenarios where target data is unavailable, providing a robust methodological approach for tackling complex tasks.
Conclusion
This study investigated the mechanisms behind the rapid learning and generalization observed during SFT. We found that attention heads are selectively activated in a task-specific manner during SFT, and that activation patterns for complex tasks can be decomposed into combinations of basic task patterns, offering a viable data substitution strategy when high-quality complex data is scarce.
Our experiments demonstrated that even small amounts of data can significantly alter activation patterns, validating the feasibility of (1) pre-training on simple tasks to equip large models with essential knowledge, and (2) fine-tuning with selected relevant data to enable rapid generalization. The improved effectiveness observed in these scenarios supports the validity of our analytical approach and offers practical solutions to data challenges in complex and specialized tasks.